# Performance of Conformal Prediction in Capturing Aleatoric Uncertainty

### v1

@@ Abstract
Conformal prediction is a model-agnostic approach to generating prediction sets that cover the true class with a high probability. Although its prediction set size is expected to capture aleatoric uncertainty, there is a lack of evidence regarding its effectiveness. The literature presents that prediction set size can upper-bound aleatoric uncertainty or that prediction sets are larger for difficult instances and smaller for easy ones, but a validation of this attribute of conformal predictors is missing. This work investigates how effectively conformal predictors quantify aleatoric uncertainty, specifically the inherent ambiguity in datasets caused by overlapping classes. We perform this by measuring the correlation between prediction set sizes and the number of distinct labels assigned by human annotators per instance. We further assess the similarity between prediction sets and human-provided annotations. We use three conformal prediction approaches to generate prediction sets for eight deep learning models trained on four datasets. The datasets contain annotations from multiple human annotators (ranging from five to fifty participants) per instance, enabling the identification of class overlap. We show that the vast majority of the conformal prediction outputs show a very weak to weak correlation with human annotations, with only a few showing moderate correlation. These findings underscore the necessity of critically reassessing the prediction sets generated using conformal predictors. While they can provide a higher coverage of the true classes, their capability in capturing aleatoric uncertainty remains limited.
@@ 1 Introduction
Quantifying model uncertainty is critical for improving model transparency and can enhance the performance of downstream tasks such as active learning and out-of-distribution (OOD) detection [20, 21]. It is usually categorized into aleatoric uncertainty (inherent to the data and irreducible) and epistemic uncertainty (related to model choice and reducible) to satisfy each task’s requirements [30, 14, 17, 8]. For example, in active learning, we would want to deprioritize samples with high aleatoric and low epistemic uncertainty [22]. Similarly, in OOD detection, we want to avoid confounding previously seen (in-distribution) high aleatoric uncertainty samples with high epistemic uncertainty samples.
@@
CP captures aleatoric uncertainty by calibrating prediction sets whose size adapts to the ambiguity in the data, regardless of the underlying model architecture [28, 7, 6]. This ability stems from using a conformity score, often derived from model outputs such as softmax probabilities, whose entropy quantifies aleatoric uncertainty on a per-instance basis [22]. Given class probabilities that are log-concave and that achieve near-perfect calibration, we can assume that the prediction set size, $|\mathcal{C}(X)|$, and an existing aleatoric uncertainty quantification method, softmax entropy [22], are positively correlated. We show evidence of this in the later parts of the paper. Correia et al. [9] also provide theoretical evidence that the prediction set size upper bounds the conditional entropy $H(X|Y)$, which captures the true intrinsic uncertainty in the distribution of labels $Y$ and inputs $X$. However, there is no clear empirical evidence of how good conformal predictors are at capturing aleatoric uncertainty, or if they do at all. In this work, we aim to address this gap.
Beyond theoretical interest, understanding what prediction set sizes represent is essential for their effective use in downstream tasks, particularly in decision-support systems that involve human interaction. User studies have shown that CP can improve human-AI complementarity in automated decision-support systems [33, 32, 10]. In such experiments, the size of the prediction sets is assumed to determine the difficulty of a classification task for human participants. In this work, we aim to evaluate this assumption empirically.
Aleatoric uncertainty arises from inherent variability that cannot be resolved through additional data or model refinement [15, 14]. In this work, we specifically focus on aleatoric uncertainty stemming from class overlap, i.e., cases where semantically distinct classes exhibit similar visual or structural characteristics in the input space, rendering accurate classification inherently uncertain. This form of uncertainty is often revealed through human annotation studies, where multiple annotators are tasked with labeling the same instance. When annotators consistently assign multiple distinct labels to a single instance, either individually or across annotators, this pattern suggests the presence of class overlap. It indicates that the resulting uncertainty is not due to noise or labeling error, but rather reflects intrinsic ambiguity in the data itself [19, 35].
@@
In this paper, we focus on datasets that contain multiple human annotations per instance. This enables us to identify and quantify class overlap as a source of aleatoric uncertainty. We then investigate whether this human-perceived ambiguity correlates with the uncertainty captured by conformal predictors, as reflected in the size of their prediction sets. While annotator disagreement has also been used as a proxy for aleatoric uncertainty [34, 16, 21], we focus on class overlap since we are interested in assessing the prediction sets. We also evaluate the performance of these prediction sets and analyze how their size influences both coverage and the strength of the correlation with class overlap. Additionally, we employ similarity metrics to assess if the prediction sets correctly identify labels collected from the multiple annotators. We define an instance as exhibiting class overlap if it contains two or more distinct class labels from an individual or multiple annotators. The degree of class overlap is quantified by the number of distinct labels assigned to the instance.
@@
  • We show that the employed conformal predictors frequently construct prediction sets that correlate weakly with the dataset’s inherent class overlap.
@@ 2 Related work
While ambiguity in human annotations has been proposed as a benchmark for model uncertainty, conformal predictors are typically assessed solely based on their coverage guarantees, i.e., whether the true class falls within the prediction set [2, 7], without regard to whether their prediction sets align with regions of genuine ambiguity in the data. Although coverage is central to the CP framework, it does not capture the informativeness of the prediction sets themselves. Despite claims that prediction set size reflects aleatoric uncertainty, this assumption remains largely untested. There is a lack of rigorous evaluations that assess whether variation in prediction set size corresponds meaningfully to inherent data ambiguity.
@@ 4.1 Datasets
For each dataset, we searched for the best target error value $\alpha\in\{0.05,0.1,0.15,0.2\}$ that minimizes the ratio of the CP mean set size to coverage. This ensures high coverage while keeping the prediction sets concise and informative. Apart from the ImageNet-ReaL dataset, which resulted in an $\alpha=0.10$, the rest of the datasets led to $\alpha=0.05$.
@@ 4.2 Models
We used eight deep learning architecture: ResNet18, ResNet34, ResNet50, Vgg16, Vgg19, DenseNet121, DenseNet161, and MobileNet-v2.111https://pytorch.org/vision/main/models.html All models were trained from scratch using an Adam optimizer with a decaying learning rate of 1e-4 and a categorical cross-entropy loss.
@@ 4.3 Evaluation
We use Spearman’s rank correlation coefficient ($r_{s}$) to assess the correlation between the prediction set size and the class overlap or the number of distinct human labels. We assess the similarity between the prediction sets and human annotations using the similarity metrics precision, recall, subset-accuracy, and hamming loss. Instances with empty prediction sets are excluded from the correlation and similarity analysis.
Table 1: Expected Calibration Error (ECE) with M = 15 bins.
Models CIFAR-10H MLRSNet FER+ ImageNet-ReaL
ResNet18 0.021 0.013 0.032 0.021
ResNet34 0.026 0.023 0.035 0.027
ResNet50 0.023 0.030 0.026 0.032
Vgg16 0.016 0.100 0.037 0.020
Vgg19 0.022 0.119 0.053 0.020
DenseNet121 0.020 0.018 0.034 0.020
DenseNet161 0.021 0.019 0.038 0.047
MobileNet-v2 0.025 0.012 0.027 0.020
@@ 5 Results
5.1 Expected Calibration Error
Table 1 presents the ECE using M=15 bins for all the models. The majority of models exhibit strong calibration with a median ECE of 0.024 across all experiments. The Vgg16 and Vgg19 models show the highest degree of miscalibration on the MLRSNet dataset.
5.2 Conformal predictors’ performance
The models’ and conformal predictors’ performance on all the four datasets is shown in Tables 2-5. We report the models’ accuracy and the conformal predictors’ coverage, SSC, and mean set size ($\bar{w}$). The APS and RAPS methods consistently outperform the LAC method in covering the true class at the cost of a larger mean set size.
Notably, Vgg16 and Vgg19, which show high miscalibration on the MLRSNet dataset, also achieve lower accuracy (Table 3). Although their coverage under the LAC method is comparable to that of other models, they are consistently outperformed when using the APS and RAPS methods.
While this trend is less apparent on datasets with similar model accuracies, we observe that lower-accuracy models consistently yield larger mean prediction set sizes across all methods. This pattern is evident for VGG16, VGG19, and MobileNet-v2 on the MLRSNet dataset (Table 3), and similarly for ResNet18, VGG16, VGG19, and MobileNet-v2 on the ImageNet-ReaL dataset (Table 5).
LAC frequently generates smaller set sizes. This can be seen in Tables 2-5. LAC can also lead to empty prediction sets, as is especially the case for the models trained on the CIFAR-10H, MLRSNet, and ImageNet-Real datasets (See the distribution of the prediction set sizes in Figure A1 in the Appendix).
Conformal prediction sets constructed with the ResNet18 model for sample images from all datasets are shown in Figure 2. As shown in the figure, while all approaches successfully cover the true class, their performance at capturing the human annotation ambiguity varies and is usually limited.
Table 2: Models’ and conformal predictors’ performance on the CIFAR-10H dataset.
Models Accuracy LAC APS RAPS
Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$
ResNet18 0.930 0.944 0.894 1.040 0.977 0.953 1.357 0.978 0.952 1.361
ResNet34 0.933 0.953 0.833 1.062 0.979 0.953 1.369 0.976 0.950 1.304
ResNet50 0.936 0.953 0.500 1.050 0.974 0.939 1.263 0.975 0.959 1.285
Vgg16 0.941 0.955 0.853 1.049 0.982 0.930 1.328 0.979 0.882 1.263
Vgg19 0.939 0.951 0.000 1.039 0.976 0.907 1.264 0.973 0.895 1.207
Table 3: Models’ and conformal predictors’ performance on the MLRSNet dataset.
Models Accuracy LAC APS RAPS
Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$
ResNet18 0.927 0.951 0.000 1.081 0.984 0.889 1.575 0.985 0.900 1.624
ResNet34 0.923 0.951 0.831 1.090 0.982 0.500 1.500 0.982 0.750 1.526
ResNet50 0.918 0.949 0.780 1.113 0.978 0.667 1.504 0.978 0.667 1.508
Vgg16 0.872 0.950 0.750 1.613 0.955 0.833 1.798 0.963 0.887 2.059
Vgg19 0.844 0.955 0.889 2.106 0.958 0.875 2.352 0.974 0.953 3.386
Table 4: Models’ and conformal predictors’ performance on the FER+ dataset.
Models Accuracy LAC APS RAPS
Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$
ResNet18 0.825 0.941 0.818 1.561 0.961 0.920 2.250 0.960 0.920 2.233
ResNet34 0.839 0.939 0.934 1.484 0.958 0.936 2.085 0.958 0.933 2.138
ResNet50 0.838 0.940 0.857 1.493 0.959 0.800 2.027 0.961 0.857 2.148
Vgg16 0.835 0.946 0.875 1.530 0.970 0.952 2.274 0.970 0.951 2.318
Vgg19 0.823 0.943 0.935 1.561 0.964 0.875 2.222 0.966 0.923 2.316
Table 5: Models’ and conformal predictors’ performance on the ImageNet-ReaL dataset.
Models Accuracy LAC APS RAPS
Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$ Coverage SSC $\bar{w}$
ResNet34 0.759 0.899 0.681 2.157 0.940 0.000 8.894 0.940 0.500 8.943
ResNet50 0.785 0.903 0.000 1.760 0.946 0.000 7.768 0.946 0.000 7.892
Vgg16 0.745 0.902 0.000 2.461 0.943 0.000 10.690 0.943 0.000 10.587
Vgg19 0.751 0.903 0.000 2.355 0.946 0.000 10.764 0.946 0.000 10.692
5.3 Coverage at different prediction set sizes
Larger prediction sets do not always lead to improved coverage (We show how coverage changes with increasing set sizes in Figure A2 in the Appendix). This is especially the case for the datasets with a large number of categories, MLRSNet (46 categories) and ImageNet-Real (1000 categories). However, some of the models trained on the CIFAR-10H and FER+ datasets led to better coverage with increasing set size.
@@ 5.4 Correlation analysis
An increased prevalence of larger prediction sets is associated with an improved Spearman’s $r_{s}$ with class overlap, suggesting that prediction set size positively contributes to capturing instance-level ambiguity (See Figure A3 in the Appendix).
@@ 5.5 Similarity analysis
All methods achieve low subset-accuracy and relatively higher Hamming loss, showing the substantial difference between prediction sets and human annotations. Precision and recall are influenced by prediction set size and the number of distinct human labels. Smaller prediction sets and more diverse human annotations, such as CIFAR-10H, MLRSNet, and FER+, yield higher precision. In contrast, ImageNet-ReaL exhibits lower precision and higher recall due to fewer human labels (five annotators chose from eight labels). Tabular results of our similarity analysis are presented in Tables A1, A2, A3 and A4 in the Appendix.

### v2

@@ Abstract
Conformal prediction is a model-agnostic approach to generating prediction sets that cover the true class with a high probability. Although its prediction set size is expected to capture aleatoric uncertainty, there is a lack of evidence regarding its effectiveness. The literature presents that prediction set size can upper-bound aleatoric uncertainty or that prediction sets are larger for difficult instances and smaller for easy ones, but a validation of this attribute of conformal predictors is missing. This work investigates how effectively conformal predictors quantify aleatoric uncertainty, specifically the inherent ambiguity in datasets caused by overlapping classes. We perform this by measuring the correlation between prediction set sizes and the number of distinct labels assigned by human annotators per instance. We further assess the similarity between prediction sets and human-provided annotations. We use three conformal prediction approaches to generate prediction sets for eight deep learning models trained on four datasets. The datasets contain annotations from multiple human annotators (ranging from five to fifty participants) per instance, enabling the identification of class overlap. We show that the vast majority of the conformal prediction outputs show a very weak to weak correlation with human annotations, with only a few showing moderate correlation. These findings underscore the necessity of critically reassessing the prediction sets generated using conformal predictors. While they can provide a higher coverage of the true classes, their capability in capturing aleatoric uncertainty and generating sets that align with human annotations remains limited.
@@ 1 Introduction
Quantifying model uncertainty is critical for improving model transparency and can enhance the performance of downstream tasks such as active learning and out-of-distribution (OOD) detection [21, 22]. It is usually categorized into aleatoric uncertainty (inherent to the data and irreducible) and epistemic uncertainty (related to model choice and reducible) to satisfy each task’s requirements [31, 16, 15, 18, 8]. For example, in active learning, we would want to deprioritize samples with high aleatoric and low epistemic uncertainty [23]. Similarly, in OOD detection, we want to avoid confounding previously seen (in-distribution) high aleatoric uncertainty samples with high epistemic uncertainty samples.
@@
CP captures aleatoric uncertainty by calibrating prediction sets whose size adapts to the ambiguity in the data, regardless of the underlying model architecture [29, 7, 6]. This ability stems from using a conformity score, often derived from model outputs such as softmax probabilities, whose entropy quantifies aleatoric uncertainty on a per-instance basis [23]. Softmax outputs can be overconfident and poorly aligned with true likelihoods, which can degrade the quality of CP sets [14]. Consequently, assessing their calibration, typically via metrics such as the Expected Calibration Error (ECE), is essential before using them for CP. Given class probabilities that are log-concave and that achieve near-perfect calibration, we can assume that the prediction set size, $|\mathcal{C}(X)|$, and an existing aleatoric uncertainty quantification method, softmax entropy [23], are positively correlated. We show evidence of this in the later parts of the paper. Correia et al. [9] also provide theoretical evidence that the prediction set size upper bounds the conditional entropy $H(X|Y)$, which captures the true intrinsic uncertainty in the distribution of labels $Y$ and inputs $X$. However, there is no clear empirical evidence of how good conformal predictors are at capturing aleatoric uncertainty, or if they do at all.
Beyond theoretical interest, understanding what prediction set sizes represent is essential for their effective use in downstream tasks, particularly in decision-support systems that involve human interaction. User studies have shown that CP can improve human-AI complementarity in automated decision-support systems [34, 33, 10]. In such experiments, the size of the prediction sets is assumed to determine the difficulty of a classification task for human participants. This paper evaluates this assumption empirically.
In this work, we specifically focus on aleatoric uncertainty stemming from class overlap, i.e., cases where semantically distinct classes exhibit similar visual or structural characteristics in the input space, rendering accurate classification inherently uncertain. This form of uncertainty is often revealed through human annotation studies, where multiple annotators are tasked with labeling the same instance. When annotators consistently assign multiple distinct labels to a single instance, either individually or across annotators, this pattern suggests the presence of class overlap. It indicates that the resulting uncertainty is not due to labeling error, but rather reflects intrinsic ambiguity in the data itself [20, 36].
@@
In this paper, we focus on datasets that contain multiple human annotations per instance. This enables us to identify and quantify class overlap as a source of aleatoric uncertainty. In addition to class overlap, we also consider the distribution of human annotations whenever such information is available. We then investigate whether the human-perceived ambiguity correlates with the uncertainty captured by conformal predictors, as reflected in the size of their prediction sets. While annotator disagreement has also been used as a proxy for aleatoric uncertainty [35, 17, 22], we focus on class overlap since we are interested in assessing the prediction sets. We also evaluate the performance of these prediction sets and analyze how their size influences both coverage and the strength of the correlation with class overlap. Additionally, we employ similarity metrics to assess if the prediction sets correctly identify labels collected from the multiple annotators. We define an instance as exhibiting class overlap if it contains two or more distinct class labels from an individual or multiple annotators. The degree of class overlap is quantified by the number of distinct labels assigned to the instance.
@@
  • We show that the employed conformal predictors frequently construct prediction sets that correlate weakly with the datasets’ inherent class overlap.
@@ 2 Related work
While ambiguity in human annotations has been proposed as a benchmark for model uncertainty, conformal predictors are typically assessed solely based on their coverage guarantees, i.e., whether the true class falls within the prediction set [2, 7], without regard to whether their prediction sets align with regions of genuine ambiguity in the data. Although coverage is central to the CP framework, it does not capture the informativeness of the prediction sets themselves.
@@ 4 Experiments
This section presents a description of the datasets employed in this paper, the trained models, and the evaluation methods.
@@ 4.1 Datasets
Per instance human annotation distribution is available for the CIFAR-10H and FER+ datasets. The rest of the datasets, i.e., MLRSNet and ImageNet-ReaL, only contain the distinct human-provided labels.
For each dataset, we searched for the best target error value $\alpha\in\{0.05,0.1,0.15,0.2\}$ that minimizes the ratio of the CP mean set size to coverage. This ensures high coverage while keeping the prediction sets concise and informative. Apart from the ImageNet-ReaL dataset, which resulted in an $\alpha=0.10$, the rest of the datasets led to $\alpha=0.05$. In addition, the optimal calibration set size and thus the calibration/test split is determined by progressively enlarging the calibration set to improve coverage until it saturates.
@@ 4.2 Models
We used eight deep learning architecture: ResNet18, ResNet34, ResNet50, VGG-16, VGG-19, DenseNet121, DenseNet161, and MobileNet-v2.111https://pytorch.org/vision/main/models.html All models were trained from scratch using an Adam optimizer with a decaying learning rate of 1e-4 and a categorical cross-entropy loss.
@@ 4.3 Evaluation
We use Spearman’s rank correlation coefficient ($r_{s}$) to assess the correlation between the prediction set size and the class overlap or the number of distinct human labels. In addition, we compute Spearman’s rank correlation between the prediction set size and human annotation distribution, using label entropy, in cases where label distribution information is available, i.e., for the CIFAR-10H and FER+ datasets. We assess the similarity between the prediction sets and human annotations using the similarity metrics precision, recall, subset-accuracy, and Hamming loss. Instances with empty prediction sets are excluded from the correlation and similarity analysis.
@@ 5 Results
This section contains the results of our correlation and similarity analysis and the performance of the conformal prediction methods. Table 1 and Table 2 present the correlation analysis results. Tables 3-6 contain the performance of the conformal predictors.
@@ 5.1 Correlation analysis
An increased prevalence of larger prediction sets is associated with an improved Spearman’s $r_{s}$ with class overlap, suggesting that prediction set size positively contributes to capturing instance-level ambiguity (See Figure A3 in Supplementary material).
Although restricted to the CIFAR-10H and FER+ datasets, we also examine correlations using the label entropy in human annotations, as these are the only datasets that provide annotation distribution information. The Spearman’s $r_{s}$ between prediction set size and human annotation entropy is reported in Table A9 in the Supplementary Material. Similar to the correlations against class overlap, the results are generally weak to very weak, with some cases showing moderate correlation.
@@ 5.2 Similarity analysis
All methods achieve low subset-accuracy and relatively higher Hamming loss, showing the substantial difference between prediction sets and human annotations. Precision and recall are influenced by prediction set size and the number of distinct human labels. Smaller prediction sets and more diverse human annotations, such as CIFAR-10H, MLRSNet, and FER+, yield higher precision. In contrast, ImageNet-ReaL exhibits lower precision and higher recall due to fewer human labels (five annotators chose from eight labels). Tabular results of our similarity analysis are presented in Tables A5, A6, A7 and A8 in the Supplementary material.
5.3 Expected Calibration Error
ECE of all the models is presented in Table A10 in Supplementary material. The majority of models exhibit strong calibration with a median ECE of 0.024 across all experiments. The VGG-16 and VGG-19 models show the highest degree of miscalibration on the MLRSNet dataset.
5.4 Conformal predictors’ performance
We report the models’ accuracy and the conformal predictors’ coverage, SSC, and mean set size ($\bar{w}$). The APS and RAPS methods consistently outperform the LAC method in covering the true class at the cost of a larger mean set size. Accuracy and coverage of all the models and conformal predictors on the CIFAR-10H dataset are shown in Table 3. A summary of coverage, SSC, and $\bar{w}$ for all the models is reported in Tables 4-6. Detailed tabular report of the models’ and conformal predictors’ performance on all four datasets can be found in Tables A1-A4 in Supplementary material.
Notably, VGG-16 and VGG-19, which show high miscalibration on the MLRSNet dataset, also achieve lower accuracy (See Table A2 in Supplementary material). Although their coverage under the LAC method is comparable to that of other models, they are consistently outperformed when using the APS and RAPS methods.
While this trend is less apparent on datasets with similar model accuracies, we observe that lower-accuracy models consistently yield larger mean prediction set sizes across all methods. This pattern is evident for VGG-16, VGG-19, and MobileNet-v2 on the MLRSNet dataset (See Table A2 in Supplementary material), and similarly for ResNet18, VGG-16, VGG-19, and MobileNet-v2 on the ImageNet-ReaL dataset (See Table A4 in Supplementary material).
LAC frequently generates smaller set sizes. This can be seen summarized in Table 6 (See details in Tables A1-A4 in Supplementary material). LAC can also lead to empty prediction sets, as is especially the case for the models trained on the CIFAR-10H, MLRSNet, and ImageNet-Real datasets (See the distribution of the prediction set sizes in Figure A1 in the Supplementary material).
Conformal prediction sets constructed with the ResNet18 model for sample images from all datasets are shown in Figure 2. As shown in the figure, while all approaches successfully cover the true class, their performance at capturing the ambiguity in human annotations varies and is usually limited.
Table 3: Coverage of conformal predictors’ on the CIFAR-10H dataset.
Models Accuracy Coverage
LAC APS RAPS
ResNet18 0.930 0.944 0.977 0.978
ResNet34 0.933 0.953 0.979 0.976
ResNet50 0.936 0.953 0.974 0.975
VGG-16 0.941 0.955 0.982 0.979
VGG-19 0.939 0.951 0.976 0.973
DenseNet121 0.940 0.942 0.979 0.975
DenseNet161 0.940 0.944 0.977 0.976
MobileNet-v2 0.939 0.951 0.984 0.982
Table 4: Mean ($\pm$ std) of models’ coverage for each dataset.
Dataset Accuracy Coverage
LAC APS RAPS
CIFAR-10H 0.937 $\pm$ 0.004 0.949 $\pm$ 0.005 0.978 $\pm$ 0.003 0.977 $\pm$ 0.003
MLRSNet 0.903 $\pm$ 0.038 0.952 $\pm$ 0.002 0.976 $\pm$ 0.012 0.979 $\pm$ 0.007
FER+ 0.832 $\pm$ 0.006 0.943 $\pm$ 0.002 0.963 $\pm$ 0.004 0.964 $\pm$ 0.004
ImageNet-ReaL 0.759 $\pm$ 0.023 0.902 $\pm$ 0.001 0.943 $\pm$ 0.003 0.943 $\pm$ 0.003
Table 5: Mean ($\pm$ std) of models’ SSC for each dataset.
Dataset Accuracy SSC
LAC APS RAPS
CIFAR-10H 0.937 $\pm$ 0.004 0.385 $\pm$ 0.428 0.939 $\pm$ 0.016 0.935 $\pm$ 0.031
MLRSNet 0.903 $\pm$ 0.038 0.500 $\pm$ 0.416 0.704 $\pm$ 0.321 0.810 $\pm$ 0.158
FER+ 0.832 $\pm$ 0.006 0.891 $\pm$ 0.051 0.900 $\pm$ 0.058 0.915 $\pm$ 0.038
ImageNet-ReaL 0.759 $\pm$ 0.023 0.234 $\pm$ 0.327 0.000 $\pm$ 0.000 0.062 $\pm$ 0.177
Table 6: Mean ($\pm$ std) of models’ mean width $\bar{w}$ for each dataset.
Dataset $\bar{w}$
LAC APS RAPS
CIFAR-10H 1.036 $\pm$ 0.019 1.344 $\pm$ 0.058 1.297 $\pm$ 0.053
MLRSNet 1.340 $\pm$ 0.398 1.770 $\pm$ 0.459 1.978 $\pm$ 0.750
FER+ 1.526 $\pm$ 0.033 2.188 $\pm$ 0.086 2.252 $\pm$ 0.073
ImageNet-ReaL 2.236 $\pm$ 0.438 9.379 $\pm$ 2.035 9.412 $\pm$ 2.016
5.5 Coverage at different prediction set sizes
Larger prediction sets do not always lead to improved coverage (We show how coverage changes with increasing set sizes for all the datasets and CP methods in Figure A2 in Supplementary material). This is especially the case for the datasets with a large number of categories, MLRSNet (46 categories) and ImageNet-Real (1000 categories). However, some of the models trained on the CIFAR-10H and FER+ datasets led to better coverage with increasing set size.